OpenStudy (anonymous):

I need help with this problem... I know the formual is y = (ax+b)/[x(x+5)] Vertical asymptotes: x=-5, x= 1 Horizontal asymptotes: y=0 x-intercept: (-3,0), y-intercept: (0,-(3/5))

OpenStudy (amistre64):

there is no vert asymp at x=1; -5 and 0 yes

OpenStudy (amistre64):

when x=0 we get -b/0 = -3/5 aint possible

OpenStudy (amistre64):

well, b/0, but same perdicament

OpenStudy (anonymous):

Heres the answer choices... I just need it explained a. f(x) = x-3/ x^2+4x-5 b. f(x) = x+3/x^2+4x-5 c.x+3/ x^2-4x-5 d. f(x)= x-5/x^2+4x-5

OpenStudy (amistre64):

i did explain it ; there are no possible results for a and b that can be parsed from teh given information

OpenStudy (phi):

Perhaps there is a typo in the question. The answer choices imply that the denominator is (x-1)(x+5) not x(x+5)

OpenStudy (amistre64):

fuzzy logic, aint got no time for fuzzies :)

OpenStudy (phi):

(x-1)(x+5) gives asymptotes of 1 and -5

OpenStudy (anonymous):

I understand that part.. I just looked it up.. I just don't know how to the top part.

OpenStudy (amistre64):

when you make x = 0, assuming the typo that is, the botom goes to -5 b/-5 = 3/-5 compare parts to solve for b

OpenStudy (phi):

Here is your problem \[ \frac{ax+b}{(x-1)(x+5) }\] the question tells you y-intercept: (0,-(3/5)) what amistre did was put 0 in for x. you have to get -3/5 out.

OpenStudy (phi):

Can you replace the x with 0 in the fraction, and simplify it?

OpenStudy (anonymous):

No

OpenStudy (phi):

Do it step by step. The top part is (ax+b) if you put in 0 instead of x you get (a*0+b) can you simplify this?

OpenStudy (anonymous):

well, anything times 0 is zero. so, don't you take out the a because it wouldn't do anything if I found a or not

OpenStudy (phi):

yes you get 0+b or just b now do the bottom (x-1)(x+5) put in 0 for x, and simplify

OpenStudy (anonymous):

b/-5

OpenStudy (phi):

yes (0-1)(0+5) = -1*5 = -5 so b/-5 now remember y-intercept: (0,-(3/5)) that is short hand for when x=0 the fraction must equal -3/5 we found the fraction is \[ \frac{b}{-5}= \frac{-3}{5} \] Can you solve for b? You multiply both sides of the equation by -5

OpenStudy (anonymous):

cross multiply ?

OpenStudy (phi):

Cross multiply gives 5b = -5*-3 or 5b= 15 now you would multiply by 1/5 on both sides. But I would just multiply the original problem by -5 on both sides.